# AI Twin: Enhancing ESL Speaking Practice through AI Self-Clones of a Better Me

### v1

@@ Abstract.
$c$s: Human-centered computing Interactive systems and toolsccs: Human-centered computing Empirical studies in HCIccs: Applied computing Interactive learning environments
@@ 1. Introduction
To realize this vision, we introduce AI Twin (Fig. 1), a personalized self-clone—a “better me” that speaks more fluent English—designed to enhance learner motivation and engagement through supportive, in-conversation language learning. In this setting, the learner converses with an AI interlocutor, while AI Twin operates in between through in-conversation rephrasing, rendering the learner’s utterances in a more fluent and confident form spoken in the learner’s cloned voice before passing them on. In doing so, AI Twin embodies the learner’s aspirational self-image by speaking as a more proficient version of them, aiming to support learner engagement. By rephrasing rather than explicitly correcting, it maintains conversational immersion while reducing performance pressure, creating a more supportive and motivating interaction environment. We view this focus on engagement as pedagogically meaningful, as SLA and online learning research consistently link to more effective language development. Accordingly, our study addresses two research questions:
@@
    We present AI Twin, an AI self-clone that embodies learner’s Ideal L2 Self by delivering implicit, in-conversation rephrasing in the learner’s own voice.
@@
    We empirically evaluate AI Twin, showing that in-conversation rephrasing fosters greater engagement than explicit feedback, and that personalization through self-representation further amplifies this effect.
@@ 2.2. Corrective Feedback and Engagement in SLA
These pedagogical insights are increasingly relevant in AI-supported language learning environments. Conversational agents and chatbots now provide opportunities for tailored, automated feedback during or after interactions (Li et al., 2022). Recent works have integrated SLA-informed CF strategies into such systems (Kim, 2024; Chhabriya et al., 2024), and notably, recasts have been shown to remain effective in AI-mediated contexts, suggesting that implicit CF strategies may also function effectively in these environments (Kim, 2024). However, most existing AI-supported ESL systems still rely heavily on explicit error correction, which can interrupt dialogue and discourage risk-taking (Zou et al., 2023; Ericsson and Johansson, 2023). Others explore alternative architectures, such as separating conversational and feedback-providing agents in role-playing scenarios (Liang et al., 2023; Yu et al., 2023). Together, these systems illustrate the growing interest in embedding pedagogically grounded feedback mechanisms into AI-mediated learning.
@@ 3.3. System Architecture and Implementation
Whereas Section 3.2 outlined the learner-facing interaction flow, here we describe the system architecture that enables this experience. The overall pipeline is depicted in Fig. 2, and consists of an initialization step followed by a repeating conversational cycle.
@@ 3.3.2. Conversational Cycle
The AI interlocutor then generates its response using gpt-4.1-mini (OpenAI (OpenAI, )) based on the reformulated utterance, rather than the learner’s original transcription. By grounding its output in the fluent version of the input, the system ensures that subsequent dialogue progresses smoothly and coherently, without being disrupted by grammatical errors or disfluencies. The interlocutor’s textual reply is then rendered into natural speech through ElevenLabs (ElevenLabs, 2024) text-to-speech (TTS) module, completing the conversational loop. This spoken output is returned to the learner, allowing them to engage in a continuous, voice-based interaction that mirrors the dynamics of human-to-human conversation.
@@ 3.3.3. Prompt Design
To guide the LLM in generating rephrasing, we designed prompts that emphasized minimal correction. The model was instructed to preserve the learner’s intended meaning, apply only the necessary grammatical or lexical adjustments, and avoid rewriting content. Providing the full dialogue history further supported contextual accuracy. The rephrasing process is illustrated in Fig. 3, and specific prompt examples are provided in Appendix A.
@@ 4. Evaluation
We conducted a within-subject, mixed-method study to evaluate AI Twin against alternative feedback strategies. This design integrated quantitative measures of learner engagement with qualitative reflections, providing both systematic assessment and deeper insight into motivational effects. Because our primary goal was to understand learners’ in-the-moment engagement and affective responses, the evaluation does not aim to assess learning gains, which typically require sustained, repeated practice or longitudinal observation (DeKeyser, 2007). Consistent with Krashen and Terrell’s Natural Approach method, which suggests that natural interaction can lower learners’ affective filters (Krashen and Terrell, 1983), we did not impose a specific knowledge targets. Instead, sessions were designed to be open-ended and communication-focused, encouraging spontaneous language use and reducing emotional barriers that might distort early engagement. This allowed the influence of each feedback strategy on natural language use and self-directed engagement to be more clearly observed. Within this overall structure, the study focuses on how the three feedback strategies influence immediate engagement, perceived support, and motivational experience within these conversational tasks. In the following sections, we describe the study details.
@@ 4.1.1. Baseline systems
To situate the contribution of AI Twin, we compared it against two baseline conditions: (1) Explicit Feedback and (2) AI Proxy. These baselines were selected to distinguish the role of feedback style, contrasting rephrasing with direct correction, and the role of personalization, contrasting the learner’s own voice with a neutral synthetic voice. Examples of each condition are shown in Fig. 4.
@@ 4.1.2. Conversational Task Design
Study sessions were built around goal-oriented conversational tasks that framed each interaction as a role-play with four short action items. Keeping tasks brief helped maintain participant attention and reduce the potential for carry-over effects across conditions in this within-subject design (Greenwald, 1976). Fig. 5 illustrates the interface through which participants engaged in the conversational tasks during the study.
@@ 4.2. Participants
We recruited 20 adult South Korean ESL learners (ages 24-36, MM=31.1, SDSD=3.44) who reported an interest in improving their spoken proficiency. Participants were recruited through online channels, as well as snowball sampling, which helped reach a diverse pool of learners. All participants had completed at least secondary education and met the eligibility criteria of being able to read and speak in English within a web-based environment.
@@
To characterize learners’ language backgrounds, we collected participants’ self-reported proficiency using the CEFR Self-Assessment Grid for the Spoken Interaction category, presented in a Korean-translated format for clarity (Council of Europe, 2020). Reported levels ranged from A1 to C1, with most participants falling within the intermediate range (Fig. 6).
@@ 4.3. Study Procedure
The session concluded with a semi-structured interview guided by a fixed set of questions. The interview probed perceptions of the three feedback conditions with respect to emotional, cognitive, and behavioral engagement, directly aligning with our research questions. Participants were asked to reflect on how each condition influenced their confidence, sense of learning, and immersion in conversation. Combining questionnaire responses and interview accounts enabled triangulation, providing both broad patterns and nuanced perspectives on learner experience. Participants were compensated with a gift voucher equivalent to \$22 USD [^1] The specific currency is not reported in order to protect anonymity. for their time. All procedures, including recruitment, consent, and compensation, were reviewed and approved by the Behavioural Research Ethics Board.
@@ 4.4.1. Questionnaire Data
All items were adapted to reflect interactions with the AI system rather than traditional classroom teachers (e.g., “my language teacher” was changed to “AI system”). Participants responded on a 6-point Likert scale to encourage directional judgments and avoid neutral responding. Reverse-coded items were adjusted prior to analysis, and subscale scores were computed by averaging item responses within each engagement category. To examine differences across the three feedback conditions, we conducted repeated-measures ANOVAs for each engagement dimension, followed by post-hoc paired $t$-tests to identify specific condition-level contrasts. Moreover, to provide finer-grained insights beyond averaged subscale scores, we also conducted ANOVAs on each question to examine how responses varied across conditions. Details of the questionnaire items are provided in Table 5.1.1.
@@ 4.4.2. Interview Data
5. Findings
5.1. How Rephrasing Supported Learner Engagement
5.1.1. Emotional Engagement
We first observed the general effectiveness of rephrasing (AI Proxy and AI Twin) as a feedback strategy. Quantitative results showed that rephrasing, whether personalized or not, significantly enhanced emotional engagement compared to Explicit Feedback condition. A repeated measures ANOVA revealed a significant effect of condition ($F(2,38)=10.89$, $p<.001$). Post-hoc comparisons indicated that both rephrasing conditions yielded higher emotional engagement than explicit correction (AI Proxy vs. Explicit Feedback: $t(19)=3.73$, $p=.001$; AI Twin vs. Explicit Feedback: t(19)=3.46t(19)=-3.46, $p=.003$), with no difference between the two rephrasing formats (t(19)=0.48t(19)=-0.48, $p=.64$). This suggests that the conversational rephrasing format itself played an important role in improving emotional engagement. The overall comparison across conditions and engagement types is illustrated in Fig. 7. Detailed repeated measures ANOVA results are reported in Table 1, and post-hoc paired $t$-test results for emotional engagement are presented in Table 2.
Per-question analyses further showed that most emotional engagement items differed significantly across conditions, with the exception of several nervousness-related items (e.g., “I felt nervous when I had to speak during this session”, “I got so nervous I forgot things I know.”, “I got nervous when the AI asked questions which I haven’t prepared in advance.”). To clarify the direction of effects, paired $t$-tests on significantly different items revealed a consistent pattern: for the majority of items (Items 1, 2, 3, 6, 8, 11, and 12), scores in the Explicit Feedback condition were significantly lower than in both rephrasing conditions. These results reinforce that embedding corrective feedback within the conversational flow helps foster more positive emotional experiences. Full per-question statistics are provided in Table 5.1.1, and detailed post-hoc results for the items that showed significant condition differences are provided in Appendix C.
[…]
@@ 5.1.1. Emotional Engagement
[…]
@@
Qualitative accounts deepened this finding by illustrating why rephrasing was experienced as more emotionally supportive. Participants often described explicit correction as discouraging, with one noting that when every mistake was listed, “it really hurt my feelings. It felt like someone next to me was constantly saying, ‘No, that’s wrong”’ (P5). In contrast, rephrasing was experienced as less judgmental and more supportive of natural conversation. As P10 reflected, “In Explicit Feedback condition, I actually felt increasing pressure that I shouldn’t make mistakes, whereas in AI Proxy and AI Twin conditions, I was able to speak more naturally and confidently.” Together, these accounts suggest that by embedding corrections within the flow of conversation, rephrasing reduced anxiety and fostered confidence.
@@ 5.2. AI Twin added value in positive engagement
Although statistical tests revealed no significant differences between the two rephrasing conditions—AI Proxy and AI Twin, participants’ accounts suggested that the AI Twin provided unique benefits across emotional, cognitive, and behavioral engagement. To address whether these benefits might stem from first-encounter effects rather than the design itself, we conducted a targeted order analysis comparing participants who experienced AI Twin first ($N=11$) with those who first interacted with AI Proxy ($N=9$). Mixed ANOVAs revealed no significant Condition $\times$ Order interactions (emotional: $F(2,36)=0.84$, $p=.44$; cognitive: $F(2,36)=0.49$, $p=.61$; behavioral: $F(2,36)=1.25$, $p=.30$). While not definitive, this pattern offers no strong indication that encountering the voice-cloned system first substantially shaped the observed effects.
@@ 5.3. AI Twin enhanced motivation
“Getting feedback in my own voice gave me the strongest feeling of wanting to improve myself. […] When I heard it in my own voice, I felt this emotion of ‘I want to speak well,’ so I paid more attention to speaking with better expressions.”(P16)
@@ 5.4. Preferences Reflected a Balance of Engagement and Learning
Post-hoc descriptive analysis explored how learners’ CEFR levels may shape their experience with the systems. Among basic users (A1–A2), only 1 of 5 preferred the AI Twin; the remaining four favored Explicit Feedback ($N=2$) and AI Proxy ($N=2$). Notably, both proficient users (C1) selected the AI Twin as their preferred option. To contextualize these differences, Fig. 8 presents average emotional, cognitive, and behavioral engagement scores across proficiency levels for each feedback condition. While exploratory, these descriptive patterns suggest that learners at different proficiency levels may orient toward feedback formats differently, with higher-proficiency learners potentially appearing more receptive to the fluency-oriented interaction style of the AI Twin.
@@ 8. Conclusion
This work introduced AI Twin, a personalized self-clone that rephrases learner utterances into more fluent versions and delivers them in the learner’s own voice. Our within-subject study with 20 adult ESL learners demonstrated that this design fostered significantly higher emotional engagement than explicit correction and enhanced motivation by aligning feedback with learners’ Ideal L2 Selves. Participants described AI Twin as more enjoyable, reassuring, and immersive, and many reported that hearing a more fluent version of themselves strengthened their confidence and desire to continue learning. These findings highlight the promise of self-representative AI in supporting affective and motivational dimensions of language learning that are often overlooked by systems focused solely on accuracy.
@@ Appendix C Per-Question Post-hoc Analyses
This section provides per-question pairwise post-hoc comparisons for all questionnaire items that showed significant differences in the ANOVA. The table reports paired t-tests between the three feedback conditions, including the test statistic ($t$), corresponding p-values, and within-subjects effect sizes ($d_{z}$).
[…]
P[$\arraybackslashp#1$

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790266isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Empirical studies in HCIccs: Applied computing Interactive learning environments
@@ 1. Introduction
To realize this vision, we introduce AI Twin (Figure 1), a personalized self-clone—a “better me” that speaks more fluent English—designed to enhance learner motivation and engagement through supportive, in-conversation language learning. In this setting, the learner converses with an AI interlocutor, while AI Twin operates in between through in-conversation rephrasing, rendering the learner’s utterances in a more fluent and confident form spoken in the learner’s cloned voice before passing them on. In doing so, AI Twin embodies the learner’s aspirational self-image by speaking as a more proficient version of them, aiming to support learner engagement. By rephrasing rather than explicitly correcting, it maintains conversational immersion while reducing performance pressure, creating a more supportive and motivating interaction environment. We view this focus on engagement as pedagogically meaningful, as SLA and online learning research consistently link engagement to more effective language development. Accordingly, our study addresses two research questions:
@@
    We present AI Twin, an AI self-clone that embodies the learner’s Ideal L2 Self by delivering implicit, in-conversation rephrasing in the learner’s own voice.
@@
    We empirically evaluate AI Twin, showing in-conversation rephrasing fosters greater engagement than explicit feedback, and that personalization through self-representation further amplifies this effect.
@@ 2.2. Corrective Feedback and Engagement in SLA
These pedagogical insights are becoming increasingly relevant in AI-supported language learning environments. Conversational agents and chatbots now provide opportunities for tailored, automated feedback during or after interactions (Li et al., 2022). Recent works have integrated SLA-informed CF strategies into such systems (Kim, 2024; Chhabriya et al., 2024), and notably, recasts have been shown to remain effective in AI-mediated contexts, suggesting that implicit CF strategies may also function effectively in these environments (Kim, 2024). However, most existing AI-supported ESL systems still rely heavily on explicit error correction, which can interrupt dialogue and discourage risk-taking (Zou et al., 2023; Ericsson and Johansson, 2023). Others explore alternative architectures, such as separating conversational and feedback-providing agents in role-playing scenarios (Liang et al., 2023; Yu et al., 2023). Together, these systems illustrate the growing interest in embedding pedagogically grounded feedback mechanisms into AI-mediated learning.
@@ 3.3. System Architecture and Implementation
Whereas Section 3.2 outlined the learner-facing interaction flow, here we describe the system architecture that enables this experience. The overall pipeline is depicted in Figure 2, and consists of an initialization step followed by a repeating conversational cycle.
@@ 3.3.2. Conversational Cycle
The AI interlocutor then generates a response via gpt-4.1-mini (OpenAI (OpenAI, )) based on the reformulated utterance, rather than the learner’s original transcription. By grounding its output in the fluent version of the input, the system ensures that subsequent dialogue progresses smoothly and coherently, without being disrupted by grammatical errors or disfluencies. The interlocutor’s textual reply is then rendered into natural speech through ElevenLabs (ElevenLabs, 2024) text-to-speech (TTS) module, completing the conversational loop. This spoken output is returned to the learner, allowing them to engage in a continuous, voice-based interaction that mirrors the dynamics of human-to-human conversation.
@@ 3.3.3. Prompt Design
To guide the LLM in generating rephrasing, we designed prompts that emphasized minimal correction. The model was instructed to preserve the learner’s intended meaning, apply only the necessary grammatical or lexical adjustments, and avoid rewriting content. Providing the full dialogue history further supported contextual accuracy. The rephrasing process is illustrated in Figure 3, and specific prompt examples are provided in Appendix A.
@@ 4. Evaluation
We conducted a within-subject, mixed-method study to evaluate AI Twin against alternative feedback strategies. This design integrated quantitative measures of learner engagement with qualitative reflections, providing both systematic assessment and deeper insight into motivational effects. Because our primary goal was to understand learners’ in-the-moment engagement and affective responses, the evaluation did not aim to assess learning gains, which typically require sustained, repeated practice or longitudinal observation (DeKeyser, 2007). Consistent with Krashen and Terrell’s Natural Approach method, which suggests that natural interaction can lower learners’ affective filters (Krashen and Terrell, 1983), we did not impose a specific knowledge targets. Instead, sessions were designed to be open-ended and communication-focused, encouraging spontaneous language use and reducing emotional barriers that might distort early engagement. This allowed the influence of each feedback strategy on natural language use and self-directed engagement to be more clearly observed. Within this overall structure, the study focuses on how the three feedback strategies influence immediate engagement, perceived support, and motivational experience within these conversational tasks. In the following sections, we describe the study details.
@@ 4.1.1. Baseline systems
To situate the contribution of AI Twin, we compared it against two baseline conditions: (1) Explicit Feedback and (2) AI Proxy. These baselines were selected to distinguish the role of feedback style, contrasting rephrasing with direct correction, and the role of personalization, contrasting the learner’s own voice with a neutral synthetic voice. Examples of each condition are shown in Figure 4.
@@ 4.1.2. Conversational Task Design
Study sessions were built around goal-oriented conversational tasks that framed each interaction as a role-play with four short action items. Keeping tasks brief helped maintain participant attention and reduce the potential for carry-over effects across conditions in this within-subject design (Greenwald, 1976). Figure 5 illustrates the interface through which participants engaged in the conversational tasks during the study.
@@ 4.2. Participants
We recruited 20 adult South Korean ESL learners (ages 24–36, MM=31.1, SDSD=3.44) who reported an interest in improving their spoken proficiency. Participants were recruited through online channels, as well as snowball sampling, which helped reach a diverse pool of learners. All participants had completed at least secondary education and met the eligibility criteria of being able to read and speak in English within a web-based environment.
@@
To characterize learners’ language backgrounds, we collected participants’ self-reported proficiency using the Common European Framework of Reference for Languages (CEFR) Self-Assessment Grid for the Spoken Interaction category, presented in a Korean-translated format for clarity (Council of Europe, 2020). Reported levels ranged from A1 to C1, with most participants falling within the intermediate range (Figure 6).
@@ 4.3. Study Procedure
The session concluded with a semi-structured interview guided by a fixed set of questions. The interview probed perceptions of the three feedback conditions with respect to emotional, cognitive, and behavioral engagement, directly aligning with our research questions. Participants were asked to reflect on how each condition influenced their confidence, sense of learning, and immersion in conversation. Combining questionnaire responses and interview accounts enabled triangulation, providing both broad patterns and nuanced perspectives on learner experience. Participants were compensated with a gift voucher equivalent to 30,000 KRW (approx. 22 USD). for their time. All procedures, including recruitment, consent, and compensation, were reviewed and approved by the Behavioural Research Ethics Board.
@@ 4.4.1. Questionnaire Data
All items were adapted to reflect interactions with the AI system rather than traditional classroom teachers (e.g., “my language teacher” was changed to “AI system”). Participants responded on a 6-point Likert scale to encourage directional judgments and avoid neutral responding. Reverse-coded items were adjusted prior to analysis, and subscale scores were computed by averaging item responses within each engagement category. To examine differences across the three feedback conditions, we conducted repeated-measures ANOVAs for each engagement dimension, followed by post-hoc paired $t$-tests to identify specific condition-level contrasts. Moreover, to provide finer-grained insights beyond averaged subscale scores, we also conducted ANOVAs on each question to examine how responses varied across conditions. Details of the questionnaire items are provided in Table 3.
@@ 4.4.2. Interview Data
5. Findings
5.1. How Rephrasing Supported Learner Engagement
5.1.1. Emotional Engagement
We first observed the general effectiveness of rephrasing (AI Proxy and AI Twin) as a feedback strategy. Quantitative results showed that rephrasing, whether personalized or not, significantly enhanced emotional engagement compared to Explicit Feedback condition. A repeated measures ANOVA revealed a significant effect of condition ($F(2,38)=10.89$, $p<.001$). Post-hoc comparisons indicated that both rephrasing conditions yielded higher emotional engagement than explicit correction (AI Proxy vs. Explicit Feedback: $t(19)=3.73$, $p=.001$; AI Twin vs. Explicit Feedback: t(19)=3.46t(19)=-3.46, $p=.003$), with no difference between the two rephrasing formats (t(19)=0.48t(19)=-0.48, $p=.64$). This suggests that the conversational rephrasing format itself played an important role in improving emotional engagement. The overall comparison across conditions and engagement types is illustrated in Figure 7. Detailed repeated measures ANOVA results are reported in Table 1, and post-hoc paired $t$-test results for emotional engagement are presented in Table 2.
[…]
Per-question analyses further showed that most emotional engagement items differed significantly across conditions, with the exception of several nervousness-related items (e.g., “I felt nervous when I had to speak during this session”, “I got so nervous I forgot things I know.”, “I got nervous when the AI asked questions which I haven’t prepared in advance.”). To clarify the direction of effects, paired $t$-tests on significantly different items revealed a consistent pattern: for the majority of items (Items 1, 2, 3, 6, 8, 11, and 12), scores in the Explicit Feedback condition were significantly lower than in both rephrasing conditions. These results reinforce that embedding corrective feedback within the conversational flow helps foster more positive emotional experiences. Full per-question statistics are provided in Table 3, and detailed post-hoc results for the items that showed significant condition differences are provided in Appendix C.
@@ 5.1.1. Emotional Engagement
# Item Explicit Feedback AI Proxy AI Twin p
Emotional Engagement
@@
Qualitative accounts deepened this finding by illustrating why rephrasing was experienced as more emotionally supportive. Participants often described explicit correction as discouraging, with one noting that when every mistake was listed, “it really hurt my feelings. It felt like someone next to me was constantly saying, ‘No, that’s wrong’ ” (P5). In contrast, rephrasing was experienced as less judgmental and more supportive of natural conversation. As P10 reflected, “In Explicit Feedback condition, I actually felt increasing pressure that I shouldn’t make mistakes, whereas in AI Proxy and AI Twin conditions, I was able to speak more naturally and confidently.” Together, these accounts suggest that by embedding corrections within the flow of conversation, rephrasing reduced anxiety and fostered confidence.
@@ 5.2. AI Twin added value in positive engagement
Although statistical tests revealed no significant differences between the two rephrasing conditions—AI Proxy and AI Twin, participants’ accounts suggested that the AI Twin provided unique benefits across emotional, cognitive, and behavioral engagement. To assess whether these benefits stem from first-encounter effects rather than the design itself, we conducted a targeted order analysis comparing participants who experienced AI Twin first ($N=11$) with those who first interacted with AI Proxy ($N=9$). Mixed ANOVAs revealed no significant Condition $\times$ Order interactions (emotional: $F(2,36)=0.84$, $p=.44$; cognitive: $F(2,36)=0.49$, $p=.61$; behavioral: $F(2,36)=1.25$, $p=.30$). While not definitive, this pattern offers no strong indication that encountering the voice-cloned system first substantially shaped the observed effects.
@@ 5.3. AI Twin enhanced motivation
“Receiving feedback in my own voice gave me the strongest feeling of wanting to improve myself. […] When I heard it in my own voice, I felt this emotion of ‘I want to speak well,’ so I paid more attention to speaking with better expressions.” (P16)
@@ 5.4. Preferences Reflected a Balance of Engagement and Learning
Post-hoc descriptive analysis explored how learners’ CEFR levels may shape their experience with the systems. Among basic users (A1–A2), only 1 of 5 preferred the AI Twin; the remaining four favored Explicit Feedback ($N=2$) and AI Proxy ($N=2$). Notably, both proficient users (C1) selected the AI Twin as their preferred option. To contextualize these differences, Figure 8 presents average emotional, cognitive, and behavioral engagement scores across proficiency levels for each feedback condition. Although based on a small number of high-proficiency participants ($N=2$), these exploratory descriptive patterns point to possible differences in how learners at different proficiency levels orient toward feedback formats, with higher-proficiency learners potentially being more receptive to the fluency-oriented interaction style of the AI Twin.
@@ 8. Conclusion
This work introduced AI Twin, a personalized self-clone that reformulates learner utterances into more fluent versions and delivers them in the learner’s own voice. Our within-subject study with 20 adult ESL learners demonstrated that this design fostered significantly higher emotional engagement than explicit correction and enhanced motivation by aligning feedback with learners’ Ideal L2 Selves. Participants described AI Twin as more enjoyable, reassuring, and immersive, and many reported that hearing a more fluent version of themselves strengthened their confidence and desire to continue learning. These findings highlight the promise of self-representative AI in supporting affective and motivational dimensions of language learning that are often overlooked by systems focused solely on accuracy.
@@ Appendix C Per-Question Post-hoc Analyses
This section provides per-question pairwise post-hoc comparisons for all 10 questionnaire items that showed significant differences in the ANOVA (9 emotional engagement items and 1 behavioral engagement item). Table 5 reports paired t-tests between the three feedback conditions, including the test statistic ($t$), corresponding p-values, and within-subjects effect sizes ($d_{z}$).